OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -35 ? -89 -86 -83 -80

OpenStudy (anonymous):

\(a_1 = a_1\) \(a_2 = a_1 + d\) \(a_3 = a_2 + d = a _1 + d + d = a_1 + 2d\) \(a_n = a_1 + (n - 1)d\) (Eq. 1) Use \(a_1 = 17\) and \(a_{17} = 35\) in Eq. 1 to find d. Then use your d and \(a_1\) to find \(a_{35} \) using Eq. 1.

OpenStudy (anonymous):

I'm still kind of confused. Would you be able to walk me through it? I don't really get how to plug the numbers in. @MaxwellSmart

OpenStudy (anonymous):

Use the formula I called Eq. 1. Here it is: \(a_n = a_1 + (n - 1)d\) You are given a1 = 13 and a17 = -35. Let's write the formula for a17: \(a_{17} = a_1 + (17 - 1)d\) \(a_{17} = a_1 + 16d\) Now we plug in the values of a1 and a17: \(-35 = 13 + 16d\) Can you solve for d?

OpenStudy (anonymous):

does d= -3?

OpenStudy (anonymous):

Correct. Now that you have the correct d, use Eq. 1 again. This time write it for \(\large a_{35} \). Then use the a1 you were given and the d you found.

OpenStudy (anonymous):